$SPAD/src/input fferr.input

Timothy Daly December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991.

———-

— * —

)set break resume
)spool fferr.output
)set message test on
)set message auto off
)clear all
--S 1 of 7
pf := PF 3
--R
--R
--R   (1)  PrimeField(3)
--R                                                                 Type: Domain
--E 1
--S 2 of 7
createIrreduciblePoly(6)$FFPOLY(pf)
--R
--R
--R         6
--R   (2)  ?  + ? + 2
--R                              Type: SparseUnivariatePolynomial(PrimeField(3))
--E 2
--S 3 of 7
createNormalPoly(6)$FFPOLY(pf)
--R
--R
--R         6     5    3
--R   (3)  ?  + 2?  + ?  + 1
--R                              Type: SparseUnivariatePolynomial(PrimeField(3))
--E 3
--S 4 of 7
createPrimitivePoly(3)$FFPOLY(pf)
--R
--R
--R         3
--R   (4)  ?  + 2? + 1
--R                              Type: SparseUnivariatePolynomial(PrimeField(3))
--E 4

--S 5 of 7
createIrreduciblePoly(3)$FFPOLY(pf)
--R
--R
--R         3
--R   (5)  ?  + 2? + 1
--R                              Type: SparseUnivariatePolynomial(PrimeField(3))
--E 5
--S 6 of 7
createNormalPoly(3)$FFPOLY(pf)
--R
--R
--R         3     2
--R   (6)  ?  + 2?  + 1
--R                              Type: SparseUnivariatePolynomial(PrimeField(3))
--E 6
--S 7 of 7
createPrimitivePoly(3)$FFPOLY(pf)
--R
--R
--R         3
--R   (7)  ?  + 2? + 1
--R                              Type: SparseUnivariatePolynomial(PrimeField(3))
--E 7
)spool
)lisp (bye)

———-

References^{[1] nothing}

5